# A Cross-Scene Benchmark for Open-World Drone Active Tracking

### v1

@@ 5.1 Experimental Settings
D-VAT $35_{\pm 7}$ $0.19_{\pm 0.03}$ $37_{\pm 7}$ $0.19_{\pm 0.03}$ $34_{\pm 6}$ $0.19_{\pm 0.03}$
R-VAT $\textbf{217}_{\pm 125}$ $\textbf{0.64}_{\pm 0.32}$ $\textbf{243}_{\pm 114}$ $\textbf{0.76}_{\pm 0.26}$ $\textbf{178}_{\pm 105}$ $\textbf{0.60}_{\pm 0.26}$
Table 3: Results of cross domain on the DAT benchmark.
5.2 Comparison Experiments
Version citystreet desert village downtown lake farmland
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ CR $TSR$
Within Scene
w/o CBT $54_{\pm 7}$ $0.30_{\pm 0.05}$ $302_{\pm 99}$ $0.75_{\pm 0.32}$ $\textbf{255}_{\pm 118}$ $\textbf{0.82}_{\pm 0.16}$ $57_{\pm 15}$ $0.33_{\pm 0.06}$ $\textbf{187}_{\pm 123}$ $0.57_{\pm 0.28}$ $46_{\pm 2}$ $0.23_{\pm 0.01}$
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{0.80}_{\pm 0.30}$ $\textbf{307}_{\pm 124}$ $\textbf{0.84}_{\pm 0.29}$ $239_{\pm 134}$ $0.73_{\pm 0.32}$ $\textbf{203}_{\pm 119}$ $\textbf{0.65}_{\pm 0.30}$ $181_{\pm 116}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$
Cross Scene
w/o CBT $39_{\pm 12}$ $0.23_{\pm 0.08}$ $\textbf{243}_{\pm 112}$ $0.69_{\pm 0.28}$ $\textbf{159}_{\pm 108}$ $0.52_{\pm 0.26}$ $50_{\pm 9}$ $0.28_{\pm 0.03}$ $139_{\pm 88}$ $0.49_{\pm 0.26}$ $53_{\pm 16}$ $0.26_{\pm 0.07}$
R-VAT $\textbf{144}_{\pm 111}$ $\textbf{0.52}_{\pm 0.29}$ $230_{\pm 115}$ $\textbf{0.67}_{\pm 0.27}$ $156_{\pm 119}$ $\textbf{0.55}_{\pm 0.31}$ $\textbf{201}_{\pm 121}$ $\textbf{0.64}_{\pm 0.30}$ $\textbf{163}_{\pm 115}$ $\textbf{0.51}_{\pm 0.29}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$
Table 4: Effectiveness of Curriculum-Based Training strategy within and across scenes on the DAT benchmark.
We compare our R-VAT with the SOTA methods for within-scene performance and cross-scene cross-domain generalization performance on the DAT benchmark. As shown in Fig. 5, our method achieves consistently higher and steadily increasing rewards throughout training, demonstrating its effectiveness in improving tracking performance across scenarios compared to other methods.
Within-scene performance. We train the model on all scenes and evaluate it on the original scene. Our R-VAT performs significantly better than other methods as shown in Tab. 2. For the $CR$, the average performance improvement on six maps relative to the D-VAT method is $591\%(35\!\rightarrow\!242)$. Regarding the $TSR$, the average enhancement is $279\%(0.19\!\rightarrow\!0.72)$.
Cross-scene performance. Our method demonstrates strong cross-scene generalization, as shown in Tab. 2. Specifically, R-VAT achieves a $376\%(37\!\rightarrow\!176)$ improvement in average $CR$ and a $200\%(0.19\!\rightarrow\!0.57)$ improvement in average $TSR$ compared to D-VAT.
Cross-domain performance. As shown in Tab. 3, our method outperforms existing methods significantly in cross-domain generalization. Specifically, the R-VAT demonstrates an average $CR$ enhancement of $509\%(35\!\rightarrow\!213)$ relative to D-VAT and $TSR$ boost of $253\%(0.19\!\rightarrow\!0.67)$.
Version night foggy snow
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
w/o CBT $136_{\pm 58}$ $0.49_{\pm 0.16}$ $138_{\pm 65}$ $0.48_{\pm 0.17}$ $107_{\pm 41}$ $0.39_{\pm 0.14}$
R-VAT $\textbf{217}_{\pm 125}$ $\textbf{0.64}_{\pm 0.32}$ $\textbf{243}_{\pm 114}$ $\textbf{0.76}_{\pm 0.26}$ $\textbf{178}_{\pm 105}$ $\textbf{0.60}_{\pm 0.26}$
Table 5: Effectiveness of Curriculum-Based Training strategy across domain on the DAT benchmark.
5.3 Ablation Experiments
We conduct ablation experiments on a Curriculum-Based Training strategy to demonstrate its effectiveness in challenging scenarios. Moreover, we conduct a validation on the farmland map to assess the effectiveness of the data augmentation techniques discussed in Sec. 4.2, as well as the reward design outlined in Sec. 4.3. Detailed descriptions of these methods are provided in Appendix E.
Effectiveness of Curriculum-Based Training strategy. As shown in Tab. 4 and Tab. 5, the R-VAT method exhibits the most significant improvements over the reinforcement learning approach on maps citystreet, downtown, and farmland, which are characterized by dense elements and visual complexity. In the within-scene, cross-scene, and cross-domain tests, enhancements in the $CR$ are $61\%(150\!\rightarrow\!242)$, $54\%(114\!\rightarrow\!176)$, and $68\%(127\!\rightarrow\!213)$, respectively, while improvements in the $TSR$ are $44\%(0.50\!\rightarrow\!0.72)$, $44\%(0.41\!\rightarrow\!0.59)$, and $49\%(0.45\!\rightarrow\!0.67)$.
Effectiveness of data augmentation. We apply randomization of angles (AR), horizontal displacement (HR), vertical displacement (VR), and gimbal pitch angle (PR) to enhance the tracking ability of the agent (see Sec. 4.2 for details). As shown in Tab. 6, adjusting AR or PR significantly improves the agent’s performance, indicating that enhancing exploratory behavior through angle variations and improving spatial awareness by randomizing the gimbal pitch angle yield positive outcomes.
Effectiveness of reward design. We contrast the performance of the R-VAT method when using the reward defined in Eq. 2 and that in [18]. As shown in Tab. 7, significant performance enhancements (about 800% improvement in $TSR$ across-scene and cross-domain) are evident on the farmland map with the utilization of Eq. 2, underscoring the pronounced effectiveness of the proposed reward design in this work. More experiments can be seen in Appendix E.
Version Within-Scene Cross-Scene Cross-Domain
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
@@ 5.3 Ablation Experiments
R-VAT $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$ $\textbf{222}_{\pm 110}$ $\textbf{0.66}_{\pm 0.27}$
Table 6: Effectiveness of augmentation on the DAT benchmark.
Version Within-Scene Cross-Scene Cross-Domain
@@
$\text{R}_{\text{D-VAT}}$ ${9}_{\pm 1}$ ${0.06}_{\pm 0.00}$ ${9}_{\pm 1}$ ${0.06}_{\pm 0.01}$ ${9}_{\pm 0}$ ${0.06}_{\pm 0.00}$
R-VAT $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$ $\textbf{222}_{\pm 110}$ $\textbf{0.66}_{\pm 0.27}$
Table 7: Effectiveness of reward design on the DAT benchmark.
@@ 6 Conclusion and Potential Impacts
In this paper, we propose a unified cross-scene cross-domain benchmark for open-world drone active tracking, called DAT. The DAT benchmark provides 24 visually complex environments to assess the algorithms’ cross-scene and cross-domain generalization abilities, and high-fidelity modeling of realistic robot dynamics. Additionally, we propose a reinforcement learning-based drone tracking method called R-VAT, aiming to improve the performance of drone tracking targets in complex scenarios. Specifically, inspired by curriculum learning, we implement a Curriculum-Based Training strategy that systematically improves agent performance in increasingly challenging scenarios. We design a goal-centered reward function to deliver precise feedback to the drone agent, preventing targets farther from the center of view from receiving higher rewards than closer ones. This allows the drone agent to adapt to the diverse motion behavior of open-world targets. Experiments demonstrate that our method is significantly superior to the SOTA method.
The DAT benchmark is a high-fidelity, cross-scene cross-domain benchmark for advancing research in drone visual active tracking. It has the potential to impact several key areas, including: 1) Forgetting in Reinforcement Learning, 2) Robustness in Reinforcement Learning, 3) Multi-Agent Reinforcement Learning, and 4) Sim-to-Real Deployment.
@@ A.1 Passive Object Tracking
Most of the proposed visual tracking benchmarks belong to passive visual tracking. LaSOT [21] and OTB2015 [64] benchmarks contain a large number of ground-based videos. These benchmarks include target videos, and the tracking algorithms utilize both the video frames and the target labels for tracking. However, ground cameras tend to be affected by occlusion and suffer from the shortcoming of limited perceptual range, so the need for drone viewpoint tracking is gradually increasing in practical applications. UAV123 [37] and VisDrone2019 [19] benchmarks are proposed for drone viewpoint, expanding the spatial dimension of perception. Meanwhile, the single-object tracking benchmarks have difficulties for many targets. MOT20 [16] and TAO [15] benchmarks are proposed for multi-object tracking to solve the above problems. In addition, the above benchmarks include videos from the RGB camera. The RGB camera’s recognition capabilities are limited in complex scenes, such as ocean environments, and challenging weather conditions, including nighttime and foggy. IPATCH [41] provides extra infrared images and other sensors like GPS to supplement the information of the sea scene. Huang et al. propose Anti-UAV410 [25], which provides infrared camera images for drone tracking.
Visual object tracking methods can be categorized into three main types: Tracking by Detection, Detection and Tracking (D&T), and pure tracking. Tracking by Detection methods [6, 62, 7] treat tracking as a sequence of independent detection tasks. These methods use object detection algorithms [44, 46] to identify the target object in each frame, connecting the detections through data association methods [27, 63] for continuous tracking. While effective in multi-target tracking, these methods may suffer from high computational demands and issues with target occlusion. D&T approaches [59, 72, 42] integrate detection and tracking, creating end-to-end models that ensure seamless information flow and reduce redundant calculations through shared feature extraction networks. Pure tracking methods can be categorized into two main types: Correlation Filters (CF) [65, 38, 24] and Siamese Networks (SN) [29, 5, 56]. CF-based models train correlation filters on regions of interest, while SN-based models compare target templates with search areas to enable precise single-target tracking.
@@ A.2 Visual Active Tracking
Passive visual tracking often falls short in real-world scenarios due to the highly dynamic nature of most targets. Visual Active Tracking (VAT) aims to autonomously follow a target object by controlling the motion system of the tracker based on visual observations [34, 73, 68]. Thus, VAT offers a more practical yet challenging solution for effective tracking in dynamic environments. Maalouf et al. [34] propose a two-stage tracking method (named FAn), which is based on a tracking model and a PID control model. This method accomplishes the fusion of perception and decision-making by transferring control information from the visual tracking model to the control model. However, the visual network necessitates extensive human labeling effort and the control model requires parameter adjustments for each scene, significantly constraining the model’s generalizability. Recently, many approaches [33, 17, 73, 18] model the VAT task as a Markov Decision Process and employ end-to-end training with reinforcement learning, resulting in a significant enhancement of the agent’s generalizability.
The complexity and diversity of VAT benchmarks are crucial for training agents with high generalizability. One common approach [18, 17, 73] to enhancing environmental diversity involves modifying texture features and lighting conditions within a single scene. However, these methods often result in low scene fidelity and unrealistic object placement. While UE4 [22] is used to create photorealistic environments in some benchmarks [73, 33], these benchmarks still face limitations in diversity and map size. Furthermore, the scenarios provided by these methods are often task-specific, offering limited configurability and lacking a unified benchmark for VAT tasks.
Existing approaches to VAT frequently neglect the randomness of target trajectories and the scalability of platforms. Target trajectories are typically predefined by rule-based patterns [18, 17, 33], which significantly restrict the exploration space. Zhong et al. [73] introduce learnable agents as targets, increasing trajectory randomness but adding additional cost. Most benchmarks provide only a single category of target [18, 17, 73, 33], limiting scalability and necessitating repetitive work for environment development. Zhou et al. [74] utilize CoppeliaSim [3] to provide five categories of noncooperative space objects. However, the use of a solid black background makes it unsuitable for general VAT scenarios. In contrast, our environment supports diverse, real-world target types and offers unified, lightweight management of target behaviors, ensuring both rationality and randomness in their actions.

### v2

@@ 5.1 Experimental Settings
D-VAT $35_{\pm 7}$ $0.19_{\pm 0.03}$ $37_{\pm 7}$ $0.19_{\pm 0.03}$ $34_{\pm 6}$ $0.20_{\pm 0.03}$
Ours $\textbf{217}_{\pm 125}$ $\textbf{0.64}_{\pm 0.32}$ $\textbf{243}_{\pm 114}$ $\textbf{0.76}_{\pm 0.26}$ $\textbf{178}_{\pm 105}$ $\textbf{0.60}_{\pm 0.26}$
Table 4: Results of ablation experiments on DAT.
Method Within-Scene Cross-Scene Cross-Domain
CRCR TSRTSR CRCR TSRTSR CRCR TSRTSR
$\text{R}_{\text{D-VAT}}$ ${9}_{\pm 1}$ ${0.06}_{\pm 0.00}$ ${8}_{\pm 1}$ ${0.05}_{\pm 0.00}$ ${9}_{\pm 0}$ ${0.06}_{\pm 0.00}$
w/o CBT $46_{\pm 2}$ $0.23_{\pm 0.01}$ $53_{\pm 16}$ $0.26_{\pm 0.07}$ $46_{\pm 2}$ $0.23_{\pm 0.01}$
@@
Ours  $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{162}_{\pm 106}$ $\textbf{0.54}_{\pm 0.26}$ $\textbf{222}_{\pm 110}$ $\textbf{0.65}_{\pm 0.27}$
5.2 Comparison Experiments
We compare our GC-VAT with the SOTA methods for within-scene performance and cross-scene cross-domain generalization performance on DAT benchmark. As shown in Fig. 5, our method achieves consistently higher and steadily increasing rewards throughout training, demonstrating its effectiveness. Both AOT and D-VAT methods fail to learn effective policies due to the misleading feedback from their distance-based rewards. In particular, AOT learns to quickly drive the target out of view, resulting in a rapidly declining reward curve. The results validate the theoretical analysis in Section 4.3. It is worth noting that although AOT and D-VAT exhibit low variance in their experimental results, consistently low rewards typically indicate a failure to learn effective tracking policies.
Within-scene performance. We train the model on all scenes and evaluate it on the original scene. Our GC-VAT performs significantly better than other methods as shown in Table 2. For the $CR$, the average performance improvement on six maps relative to the D-VAT method is $591\%(35\!\rightarrow\!242)$. Regarding the $TSR$, the average enhancement is $279\%(0.19\!\rightarrow\!0.72)$.
Cross-scene performance. Our method demonstrates strong cross-scene generalization, as shown in Table 2. Specifically, GC-VAT achieves a $376\%(37\!\rightarrow\!176)$ improvement in average $CR$ and a $200\%(0.19\!\rightarrow\!0.57)$ improvement in average $TSR$ compared to D-VAT.
Cross-domain performance. As shown in Table 5, our method outperforms existing methods significantly in cross-domain generalization. Specifically, GC-VAT demonstrates an average $CR$ enhancement of $509\%(35\!\rightarrow\!213)$ relative to D-VAT and $TSR$ boost of $253\%(0.19\!\rightarrow\!0.67)$.
5.3 Ablation Experiments
We conduct ablation experiments on Goal-Centered Reward to validate the results of the analysis presented in Section 4.3. Moreover, we verify whether the Curriculum-Based Training strategy and domain rondomization from Section 4.4 lead to a significant performance improvement. We present results on the farmland map in Table 5, with additional results provided in Appendix E.3.
Effectiveness of reward design. We contrast the performance of GC-VAT method when using the reward defined in Eq. 3 and that in dvat . As shown in Table 5, significant performance enhancements (about 800% improvement in $TSR$ across-scene and cross-domain) are evident with the utilization of Eq. 3. These results strongly corroborate the analysis in Section 4.3 and underscore the effectiveness of the proposed reward. See Appendix E.3 for more experimental results.
Effectiveness of CBT strategy and domain randomization. As shown in Table 5, without the CBT strategy, the model fails to learn effective tracking policies, resulting in consistently low rewards across different tests. In addition, our domain randomization approach yields significant improvements. Specifically, $AR$, HRHR, VRVR, and $PR$ denote the randomization of the drone’s initial angle, horizontal and vertical distance relative to the target, and gimbal pitch angle, respectively. Among these, $AR$ contributes the most to performance gains, indicating that encouraging diverse actions through angle randomization facilitates the agent’s exploration of optimal policies.
Robustness under wind gusts and precipitation. To further investigate the impact of real-world disturbances on the GC-VAT method, we conduct rigorous tests under wind gusts and sensor degradation caused by precipitation. Specifically, we simulate wind effects by applying randomized perturbations along the forward, lateral, and yaw directions during testing. The results are summarized in Table 6, where the model is trained on citystreet-day and evaluated on citystreet-foggy with added wind perturbations. The Tracking Success Rate (TSR) drops by less than 0.06, demonstrating that GC-VAT maintains strong robustness under significant wind disturbances. See Appendix E.3 for more details.
To simulate the blurring caused by raindrops, we follow established practices in test-time adaptation literature li2019heavy . Specifically, we train the policy on citystreet-day map and evaluate under synthetically generated rain in within-scene, cross-scene, and cross-domain settings. To ensure realism, we exclude snowy conditions from the cross-domain evaluation, as snow and rain rarely co-occur in real-world environments. The results in Table 6 show only marginal performance degradation (less than 0.07 in TSR) under rain simulation, confirming that GC-VAT is robust to blurring caused by raindrops.
Robustness to distractors and novel targets. As shown in Table 6, our model maintains high tracking performance even when a similar-looking vehicle is introduced near the target, demonstrating its ability to effectively distinguish the true target from confusers. In addition, we evaluate GC-VAT on an unseen target class (bus). As shown in Table 6, our model maintains strong tracking performance, with a TSR drop of less than 0.03 when encountering this novel object.
Table 5: Performance under wind disturbances and target distractors.
CR TSR
w/ Forward $302_{\pm 94}$ $0.91_{\pm 0.18}$
w/ Lateral $304_{\pm 82}$ $0.91_{\pm 0.19}$
w/ Yaw $301_{\pm 120}$ $0.88_{\pm 0.23}$
w/ Distractor $293_{\pm 120}$ ${0.91}_{\pm 0.15}$
Ours $\textbf{316}_{\pm 84}$ $\textbf{0.94}_{\pm 0.14}$
Table 6: Performance under rainy conditions and unseen targets. We evaluate the model trained on citystreet-day.
Method Within-Scene Cross-Scene Cross-Domain
@@ 5.3 Ablation Experiments
w/ rain $266_{\pm 110}$ $0.74_{\pm 0.29}$ $139_{\pm 109}$ $0.45_{\pm 0.30}$ $274_{\pm 103}$ $0.77_{\pm 0.29}$
Unseen Target $222_{\pm 92}$ $0.79_{\pm 0.25}$ $131_{\pm 89}$ $0.50_{\pm 0.33}$ $207_{\pm 94}$ $0.79_{\pm 0.27}$
Ours $\textbf{279}_{\pm 110}$ $\textbf{0.80}_{\pm 0.30}$ $\textbf{144}_{\pm 111}$ $\textbf{0.52}_{\pm 0.29}$ $\textbf{258}_{\pm 110}$ $0\textbf{.82}_{\pm 0.23}$
Figure 6: Results on real-world images.
Table 7: Effectiveness of GC-VAT on Sim2Real test. We select eight video sequences from each dataset for evaluation.
Video VOT VOT_TPAMI DTB70 DTB70 UAVDT UAVDT
Average Correct Action Rate
Random $0.413$ $0.426$ $0.421$
Ours 0.795 0.833 0.802
Figure 7: Schematic of the real-world deployment pipeline.
Figure 8: Results on real drones.
5.4 Experiments in Real-world Scenarios
Effectiveness on real-world images. Due to the difficulty real-robot evaluation, we follow e2e and validate GC-VAT on real images. We perform zero-shot transfer tests using 8 videos each from VOT VOT_TPAMI , DTB70 DTB70 and UAVDT UAVDT datasets. Although camera control is unavailable in recorded videos, we can feed frames into the model and verify the reasonableness of its predicted actions.
The output actions for the VOT video car16 are shown in Fig. 6. Each point represents the target position in the image, with colors indicating different actions. As Fig. 6 illustrates, when the target is located on the left (right) side, the tracker tends to move left (right), attempting to bring the target to the center. Quantitatively, we use the Correct Action Rate, i.e., the accuracy of predicted actions, to evaluate the performance. As shown in Table 6, GC-VAT achieves an average Correct Action Rate (CAR) of $81.0\%$ across 24 videos, demonstrating its effectiveness. More importantly, it is significantly superior to random policy ($p<0.001$) as verified by a t-test. See Appendix E.4 for more results.
Effectiveness on real drones. Furthermore, as a critical step beyond image-based evaluation, we conduct real-world experiments on a DJI Mini 3 Pro mini3pro drone. As shown in Fig. 8, we deploy GC-VAT on a laptop equipped with an RTX 3050 GPU and an Intel i5 CPU, use the DJI Mobile SDK msdk to obtain images, and control the drone with the predicted actions. The entire pipeline operates at over 30 FPS. As Fig. 8 illustrates, the model can output actions that maintain the target at the image center. Quantitatively, GC-VAT achieves an average zero-shot TSR of $88.4\%$ and a CAR of $81.3\%$. This successful zero-shot Sim-to-Real transfer validates the practical applicability of our approach.
@@ 6 Conclusion and Potential Impacts
In this paper, we propose the first open-world drone active air-to-ground tracking benchmark, called DAT. DAT benchmark encompasses 24 city-scale scenes, featuring targets with human-like behaviors and high-fidelity dynamics simulation. DAT also provides a digital twin tool for unlimited scene generation. DAT benchmark has the potential to impact several key areas, including: 1) Forgetting in Reinforcement Learning, 2) Robustness in Reinforcement Learning, 3) Multi-Agent Reinforcement Learning, and 4) Sim-to-Real Deployment. Additionally, we propose a reinforcement learning-based drone tracking method called GC-VAT, aiming to improve the performance of drone tracking targets in complex scenarios. Specifically, we design a Goal-Centered Reward to provide precise feedback across viewpoints to the agent, enabling it to expand perception range through unrestricted perspectives. Then we propose qualitative and theoretical methods to analyze the reward effectiveness. Moreover, inspired by curriculum learning, we implement a Curriculum-Based Training strategy that progressively improves agent performance in increasingly complex scenarios. Experiments on the simulator and real-world images validate the analysis and demonstrate that our method is significantly superior to the SOTA methods.
Acknowledgements
This work was partially supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.U24A20327).
Supplementary Materials for
“Open-World Drone Active Tracking with
Goal-Centered Rewards”
Contents
@@ A.1 Passive Object Tracking
Most of the proposed visual tracking benchmarks belong to passive visual tracking. LaSOT lasot and OTB2015 otb2015 benchmarks contain a large number of ground-based videos. These benchmarks include target videos, and the tracking algorithms utilize both the video frames and the target labels for tracking. However, ground cameras tend to be affected by occlusion and suffer from the shortcoming of limited perceptual range, so the need for drone viewpoint tracking is gradually increasing in practical applications. UAV123 UAV123 and VisDrone2019 visdrone2019 benchmarks are proposed for drone viewpoint, expanding the spatial dimension of perception. Meanwhile, the single-object tracking benchmarks have difficulties for many targets. MOT20 MOT20 and TAO TAO benchmarks are proposed for multi-object tracking to solve the above problems. In addition, the above benchmarks include videos from the RGB camera. The RGB camera’s recognition capabilities are limited in complex scenes, such as ocean environments, and challenging weather conditions, including nighttime and foggy. IPATCH IPATCH provides extra infrared images and other sensors like GPS to supplement the information of the sea scene. Huang et al. propose Anti-UAV410 antiuav410 , which provides infrared camera images for drone tracking.
Visual object tracking methods can be categorized into three main types: Tracking by Detection, Detection and Tracking (D&T), and pure tracking. Tracking by Detection methods SORT ; DeepSORT ; bi2022iemask treat tracking as a sequence of independent detection tasks. These methods use object detection algorithms yolov1 ; fasterrcnn to identify the target object in each frame, connecting the detections through data association methods hungarian ; cmatching for continuous tracking. While effective in multi-target tracking, these methods may suffer from high computational demands and issues with target occlusion. D&T approaches JDE ; FairMOT ; ChainedTracker integrate detection and tracking, creating end-to-end models that ensure seamless information flow and reduce redundant calculations through shared feature extraction networks. Pure tracking methods can be categorized into two main types: Correlation Filters (CF) CIRNET ; CA+CF ; DFCNet and Siamese Networks (SN) siamrpn ; SiamFC ; SiamMask . CF-based models train correlation filters on regions of interest, while SN-based models compare target templates with search areas to enable precise single-target tracking.
@@ A.2 Visual Active Tracking
Passive visual tracking often falls short in real-world scenarios due to the highly dynamic nature of most targets. Visual Active Tracking (VAT) aims to autonomously follow a target object by controlling the motion system of the tracker based on visual observations followanything ; advat+ ; yuan2023active . Thus, VAT offers a more practical yet challenging solution for effective tracking in dynamic environments. Maalouf et al. followanything propose a two-stage tracking method (named FAn), which is based on a tracking model and a PID control model. This method accomplishes the fusion of perception and decision-making by transferring control information from the visual tracking model to the control model. However, the visual network necessitates extensive human labeling effort and the control model requires parameter adjustments for each scene, significantly constraining the model’s generalizability. Recently, many approaches e2e ; cvat ; advat+ ; dvat model the VAT task as a Markov Decision Process and employ end-to-end training with reinforcement learning, resulting in a significant enhancement of the agent’s generalizability.
The complexity and diversity of VAT benchmarks are crucial for training agents with high generalizability. One common approach dvat ; cvat ; advat+ to enhancing environmental diversity involves modifying texture features and lighting conditions within a single scene. However, these methods often result in low scene fidelity and unrealistic object placement. While UE4 ue4 is used to create photorealistic environments in some benchmarks advat+ ; e2e , these benchmarks still face limitations in diversity and map size. Furthermore, the scenarios provided by these methods are often task-specific, offering limited configurability and lacking a unified benchmark for VAT tasks.
Existing approaches to VAT frequently neglect the randomness of target trajectories and the scalability of platforms. Target trajectories are typically predefined by rule-based patterns dvat ; cvat ; e2e , which significantly restrict the exploration space. Zhong et al. advat+ introduce learnable agents as targets, increasing trajectory randomness but adding additional cost. Most benchmarks provide only a single category of target dvat ; cvat ; advat+ ; e2e , limiting scalability and necessitating repetitive work for environment development. Zhou et al. space utilize CoppeliaSim coppeliasim to provide five categories of noncooperative space objects. However, the use of a solid black background makes it unsuitable for general VAT scenarios. In contrast, our environment supports diverse, real-world target types and offers unified, lightweight management of target behaviors, ensuring both rationality and randomness in their actions.